OpenStudy (clamin):

Which of the following functions has a range of all real numbers? A. y= -7x + 2 B. y= 4x^2 - 1 C. y= -3x^2 + 5 D. y= 6 (2^x)

OpenStudy (mathmale):

First: what does "range" mean in this context?

OpenStudy (clamin):

The subtraction of smalles and largest number

OpenStudy (clamin):

@mathmale

OpenStudy (greatlife44):

Range means the list of all possible y values.

OpenStudy (clamin):

Is it C?? @greatlife44

OpenStudy (greatlife44):

The only function that has a range of all real numbers is A. \[\mathbb{R} , all~real~numbers\] because you can get any y value that's from \[(-\infty, \infty)\]

OpenStudy (welshfella):

yes note that x^2 cannot be negative for real values of x so the last 3 options range is not all real numbers